# Singleton algorithms for the Constraint Satisfaction Problem

### v1

@@ Abstract
A natural strengthening of an algorithm for the (promise) constraint satisfaction problem is its singleton version: we first fix a constraint to some tuple from the constraint relation, then run the algorithm, and remove the tuple from the constraint if the answer is negative. We characterize the power of the singleton versions of standard universal algorithms for the (promise) CSP over a fixed template in terms of the existence of a minion homomorphism. Using the Hales-Jewett theorem, we show that for finite relational structures this minion condition is equivalent to the existence of polymorphisms with certain symmetries, called palette block symmetric polymorphisms. By proving the existence of such polymorphisms we establish that the singleton version of the BLP+AIP algorithm solves all tractable CSPs over domains of size at most 7. Finally, by providing concrete CSP templates, we illustrate the limitations of linear programming, the power of the singleton versions, and the elegance of the palette block symmetric polymorphisms.
@@ 1.1 Constraint Satisfaction Problem
In general the problem $\operatorname{FiniteCSP}$ is NP-complete, and to get tractable cases we usually restrict the set of allowed constraint. For a finite relational structure $\mathbb{A}$, also called the constraint language or template, by $\operatorname{CSP}(\mathbb{A})$ we denote the restricted version of $\operatorname{FiniteCSP}$ where each domain is $A$ and the constraint relations are from $\mathbb{A}$. Alternatively, the input $\langle X,D,\mathbf{C}\rangle$ of $\operatorname{CSP}(\mathbb{A})$ can be viewed as a relational structure $\mathbb{X}$ in the same signature $\sigma$ as $\mathbb{A}$ on the domain $X$ such that $(x_{1},\dots,x_{m})\in R^{\mathbb{X}}$ for $R\in\sigma$ if and only if $R(x_{1},\dots,x_{m})$ is a constraint from $\mathbf{C}$. Then $\operatorname{CSP}(\mathbb{A})$ is equivalent to checking whether there exists a homomorphism $\mathbb{X}\to\mathbb{A}$. Nevertheless, we prefer not to fix the target structure $\mathbb{A}$ and work with a triple $\langle X,D,\mathbf{C}\rangle$ as standard algorithms for the CSP change the constraint relations and even domains of the variables. We have the following characterization of the complexity of $\operatorname{CSP}(\mathbb{A})$ known as the CSP Dichotomy Conjecture.
@@
The NP-hardness in the above theorem follows from [4, 5] and [13]. The essential part of each proof of the CSP Dichotomy Conjecture is a polynomial algorithm that works for all tractable cases. Both polynomial algorithms [16, 15, 6, 7] we know are not universal in the sense that they work only for a fixed structure $\mathbb{A}$ and one of the main open questions in the area is the existence of a universal algorithm.
@@
1.2 Universal algorithms
@@
    if $\mathfrak{A}$ returns “No” then $\mathcal{I}$ has no solutions.
@@
  • CSP over a fixed template. The first goal is an algorithm that works correctly on all instances of $\operatorname{CSP}(\mathbb{A})$, whenever $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time (by an non-universal algorithm).
  • Perfect Matching Problem. In addition, we expect the algorithm to solve the perfect matching problem correctly.
@@
In the paper we consider classic universal algorithms such as the arc-consistency algorithm, Basic Linear Programming (BLP), Affine Integer Programming (AIP), and their combinations (see Section 3).
@@ 1.3 Main results
As shown in [2], polymorphisms $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ not only characterize the complexity of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ but also determine which algorithms can solve it. For instance, the arc-consistency algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities. Similarly, BLP and AIP are characterized by symmetric and alternating polymorphisms, respectively.
Following the ideas from [2], we characterize the power of the singleton versions of such algorithms by the existence of a polymorphism in $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ that is symmetric enough. The first main result of the paper is presented in Table 1. For every algorithm or combination of algorithms from Section 3 we provide a criterion for the algorithm to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. This criterion is formulated in terms of palette block symmetric polymorphisms, which are formally defined in Section 3.4. To explain the idea of such polymorphisms, let us consider the singleton version of the BLP. As shown in Table 1, this algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{m},\dots,\overline{m}}_{n})$-palette block symmetric function for all $m$ and $n$. Such a function has arity $m\cdot n$, with its arguments divided into $n$ blocks of $m$ variables each. We require the function to be symmetric within blocks (permutations of variables inside a block do not change the result), but only on tuples that additionally satisfy the palette property. A tuple is a palette tuple if any element that appears in it must completely fill one of the overlined blocks. For example, $(\overline{3},\overline{3},\overline{3})$-palette block symmetric function $f$ must satisfy the identity
@@
ArcCons $m$-ary totally symmetric function
@@
$\mbox{{ArcCons}}\wedge\mathrm{AIP}$ $(m,2n+1)$-block function whose first block is totally symmetric and the second block is alternating
@@
$\mbox{{ArcCons}}+\mathrm{AIP}$ $(2\ell+1)$-ary weakly alternating function
@@
$\mathrm{Singl}\mbox{{ArcCons}}$ $\mathrm{CSingl}\mbox{{ArcCons}}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette block totally symmetric function
@@
$\mathrm{Singl}\mathrm{BLP}$ $\mathrm{CSingl}\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette block symmetric function
@@
$\mathrm{CSingl}\mathrm{AIP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block alternating function
@@
$\mathrm{CSingl}(\mbox{{ArcCons}}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette block function whose odd blocks are totally symmetric and even blocks are alternating
@@
$\mathrm{Singl}(\mbox{{ArcCons}}+\mathrm{AIP})$ $\mathrm{CSingl}(\mbox{{ArcCons}}+\mathrm{AIP})$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block weakly alternating function
@@
$\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric function
@@
$\mathrm{CSingl}(\mathrm{BLP}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette block function whose odd blocks are symmetric and even blocks are alternating
@@
$\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ (CLAP) $\mathrm{CSingl}\mathrm{BLP}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette block function whose first $n$ blocks are symmetric and last block is alternating
@@
$\mathrm{Singl}\mathrm{AIP}+\mbox{{ArcCons}}$ $\mathrm{CSingl}\mathrm{AIP}+\mbox{{ArcCons}}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette block function whose first $n$ blocks are alternating and last block is totally symmetric
@@
$\mathrm{CSingl}\mathrm{AIP}+\mathrm{BLP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette block function whose first $n$ blocks are alternating and last block is symmetric
@@
$\mathrm{Singl}\mbox{{ArcCons}}+\mathrm{BLP}$ $\mathrm{CSingl}\mbox{{ArcCons}}+\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette block function whose first $n$ blocks are totally symmetric and last block is symmetric.
@@
$\mathrm{Singl}\mbox{{ArcCons}}+\mathrm{AIP}$ $\mathrm{CSingl}\mbox{{ArcCons}}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette block function whose first $n$ blocks are totally symmetric and last block is alternating.
@@
Thus, up to domain 7 (even in the multi-sorted case), all tractable CSPs can be solved by this very simple algorithm. Unfortunately, as noted by Kompatscher, the structure arising from the dihedral group $\mathbb{D}_{4}$ does not admit palette block symmetric polymorphisms, which implies that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve it.
@@
As a byproduct, we also show that the two versions of the singleton algorithms — fixing a tuple of a constraint or a value of a variable — have the same power for most of the algorithms.
@@
In Section 4 we define several CSP templates to illustrate the limitations of linear programming algorithms such as BLP, AIP, and $\mathrm{BLP}+\mathrm{AIP}$. By providing concrete unsatisfiable instances on which $\mathrm{BLP}+\mathrm{AIP}$ incorrectly returns “yes”, we demonstrate that the failure of the algorithm arises because the corresponding template is, in a certain sense, disconnected. We then show that the singleton version overcomes this difficulty, as fixing an element allows it to solve each connected part separately. Finally, we show that the template arising from the dihedral group $\mathbb{D}_{4}$ cannot be solved even by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. Thus, we identify another fundamental reason why linear programming algorithms fail. Notice that in this section we not only present instances that fool the algorithms but also demonstrate the characterization of the singleton versions of the algorithms by providing concrete palette block symmetric polymorphisms.

### v2

@@ Abstract
A natural strengthening of an algorithm for the (promise) constraint satisfaction problem is its singleton version: we first fix a constraint to a tuple from the constraint relation, then run the algorithm, and remove the tuple from the constraint if the answer is negative. We characterize the power of the singleton versions of standard universal algorithms for the (promise) CSP over a fixed template in terms of the existence of a minion homomorphism. Using the Hales-Jewett theorem, we show that for finite relational structures this minion condition is equivalent to the existence of polymorphisms with certain symmetries, which we call palette symmetric polymorphisms. By proving the existence of such polymorphisms we establish that the singleton version of the BLP+AIP algorithm solves all (multi-sorted) tractable CSPs over domains of size at most 7. We further show that already for domain size 8 there exists a relational structure arising from the dihedral group $D_{4}$ that does not admit palette symmetric polymorphisms and cannot be solved by singleton BLP+AIP. By providing concrete CSP templates, we illustrate the limitations of linear programming, the power of the singleton versions, and the elegance of palette symmetric polymorphisms. Surprisingly, we also identify a concrete temporal relational structure whose CSP can be solved by a singleton algorithm after sampling, yet it does not admit palette symmetric polymorphisms. This highlights that the Hales-Jewett argument applies only to finite structures. Nevertheless, we introduce generalized palette polymorphisms for every tractable temporal relational structure and show that they yield a rounding procedure implying that the corresponding singleton algorithm solves its CSP after sampling.
@@ 1.1 Constraint Satisfaction Problem
In general the problem $\operatorname{FiniteCSP}$ is NP-complete, and to get tractable cases we usually restrict the set of allowed constraints. For a finite relational structure $\mathbb{A}$, also called the constraint language or template, by $\operatorname{CSP}(\mathbb{A})$ we denote the restricted version of $\operatorname{FiniteCSP}$ where each domain is $A$ and the constraint relations are from $\mathbb{A}$. Alternatively, the input $\langle X,D,\mathbf{C}\rangle$ of $\operatorname{CSP}(\mathbb{A})$ can be viewed as a relational structure $\mathbb{X}$ in the same signature $\sigma$ as $\mathbb{A}$ on the domain $X$ such that $(x_{1},\dots,x_{m})\in R^{\mathbb{X}}$ for $R\in\sigma$ if and only if $R(x_{1},\dots,x_{m})$ is a constraint from $\mathbf{C}$. Then $\operatorname{CSP}(\mathbb{A})$ is equivalent to checking whether there exists a homomorphism $\mathbb{X}\to\mathbb{A}$. Nevertheless, we prefer not to fix the target structure $\mathbb{A}$ and work with a triple $\langle X,D,\mathbf{C}\rangle$ as standard algorithms for the CSP change the constraint relations and even domains of the variables. We have the following characterization of the complexity of $\operatorname{CSP}(\mathbb{A})$ known as the CSP Dichotomy Conjecture.
@@
The NP-hardness in the above theorem follows from [18, 17] and [35]. The essential part of each proof of the CSP Dichotomy Conjecture is a polynomial algorithm that works for all tractable cases. Both polynomial algorithms [43, 42, 15, 16] we know are not universal in the sense that they work only for a fixed structure $\mathbb{A}$, and one of the main open questions in the area is the existence of a universal algorithm.
@@
1.2 Universal algorithms and their singleton versions
@@
    if $\mathfrak{A}$ returns “No” on an instance $\mathcal{I}$, then $\mathcal{I}$ has no solutions.
@@
  • CSP over a fixed template. The first goal is an algorithm that works correctly on all instances of $\operatorname{CSP}(\mathbb{A})$, whenever $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time (by a non-universal algorithm).
  • Perfect Matching Problem. In addition, we expect the algorithm to solve the perfect matching problem correctly. This problem can be viewed as a CSP on $\{0,1\}$ over the relations 1IN2, 1IN3, 1IN4, … with the additional structural restriction that every variable appears exactly twice. We highlight this problem because it is a particularly elegant tractable CSP for which all known algorithms are non-universal and designed specifically for this problem.
@@
In the paper we consider classic universal algorithms such as the arc-consistency (AC) algorithm, Basic Linear Programming (BLP), Affine Integer Programming (AIP), and their combinations (see Section 3).
The main role in this paper is played by the singleton versions of the algorithms. Such algorithms also appear in the literature under the name “cohomological” [37, 34], but we prefer the term singleton, since it appeared earlier [22] and, in our opinion, better reflects the underlying idea. For a universal algorithm $\mathfrak{A}$ by $\mathrm{CSingl}\mathfrak{A}$ we denote the algorithm that gradually reduces constraint relations in the following way. It fixes a constraint to some tuple from the constraint relation, then runs the algorithm $\mathfrak{A}$, and removes the tuple from the constraint if the answer is negative. It repeats this for every constraint and every tuple until no more tuples can be removed. If some constraint relation becomes empty, it returns “No”. Otherwise, it returns “Yes”.
A slight modification of this algorithm is $\mathrm{Singl}\mathfrak{A}$, where instead of fixing a constraint to a tuple, it fixes a variable to a value and removes the value from the domain if the answer of $\mathfrak{A}$ is negative. For most of the algorithms, $\mathrm{Singl}\mathfrak{A}$ is not stronger than $\mathrm{CSingl}\mathfrak{A}$, but this is not obvious for $\mathrm{Singl}\mathrm{AIP}$, as $\mathrm{Singl}$-algorithm changes all the constraints containing a variable $x$ when removing an element from the domain of $x$ (see Remark 3.1). The precise definitions are given in Section 3.4.
There is another natural way to strengthen a universal algorithm. Instead of arc consistency, one may consider higher levels of local consistency, such as $k$-consistency [5]. One can also combine this with AIP by first enforcing some level of consistency and then applying the AIP relaxation to partial solutions [26]. Finally, one can further strengthen such algorithms by considering the singleton (cohomological) version [37]. A major open question in the area is whether some combination of a sufficient level of consistency and a sufficient level of linear relaxation can solve all tractable CSPs. Studying the limitations of such algorithms is currently a very active topic [26, 24, 34, 20]. Only recently, the seemingly simple conjecture of Dalmau and Opršal [26] that an affine relaxation above $k$-consistency can solve all tractable CSPs was disproved [34]. Several recent papers demonstrate how such algorithms can be fooled by random instances and regular graphs [20, 21, 25, 34]. Since higher levels and hierarchies are outside the scope of the present paper, we do not define them here and instead refer the reader to the above references.
@@ 1.3 Main results
As shown in [3], the polymorphisms $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ not only characterize the complexity of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ but also determine which algorithms can solve it. For instance, the arc-consistency algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities. Similarly, BLP and AIP are characterized by symmetric and alternating polymorphisms, respectively.
Following the ideas from [3], we characterize the power of the singleton versions of such algorithms by the existence of a polymorphism in $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ that is symmetric enough. The first main result of the paper is presented in Table 1. For every algorithm or combination of algorithms from Section 3, we provide a criterion for the algorithm to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. This criterion is formulated in terms of palette symmetric polymorphisms, which are formally defined in Section 3.4. Similar polymorphisms first appeared in [22] in the study of the power of the singleton arc-consistency algorithm. To explain the idea of such polymorphisms, let us consider the singleton version of the BLP. As shown in Table 1, this algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{m},\dots,\overline{m}}_{n})$-palette symmetric function for all $m$ and $n$. Such a function has arity $m\cdot n$, with its arguments divided into $n$ blocks of $m$ variables each. We require the function to be symmetric within blocks (permutations of variables inside a block do not change the result), but only on tuples that additionally satisfy the palette property. A tuple is a palette tuple if any element that appears in it must completely fill one of the overlined blocks. For example, $(\overline{3},\overline{3},\overline{3})$-palette symmetric function $f$ must satisfy the identity
@@
AC $m$-ary totally symmetric function
@@
$\mbox{{AC}}\wedge\mathrm{AIP}$ $(m,2n+1)$-block function whose first block is totally symmetric and the second block is alternating
@@
$\mbox{{AC}}+\mathrm{AIP}$ $(2\ell+1)$-ary weakly alternating function
@@
$\mathrm{Singl}\mbox{{AC}}$ $\mathrm{CSingl}\mbox{{AC}}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette totally symmetric function
@@
$\mathrm{Singl}\mathrm{BLP}$ $\mathrm{CSingl}\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette symmetric function
@@
$\mathrm{CSingl}\mathrm{AIP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette alternating function
@@
$\mathrm{CSingl}(\mbox{{AC}}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette function whose odd blocks are totally symmetric and even blocks are alternating
@@
$\mathrm{Singl}(\mbox{{AC}}+\mathrm{AIP})$ $\mathrm{CSingl}(\mbox{{AC}}+\mathrm{AIP})$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette weakly alternating function
@@
$\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric function
@@
$\mathrm{CSingl}(\mathrm{BLP}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette function whose odd blocks are symmetric and even blocks are alternating
@@
$\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ (CLAP) $\mathrm{CSingl}\mathrm{BLP}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette function whose first $n$ blocks are symmetric and last block is alternating
@@
$\mathrm{Singl}\mathrm{AIP}+\mbox{{AC}}$ $\mathrm{CSingl}\mathrm{AIP}+\mbox{{AC}}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette function whose first $n$ blocks are alternating and last block is totally symmetric
@@
$\mathrm{CSingl}\mathrm{AIP}+\mathrm{BLP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette function whose first $n$ blocks are alternating and last block is symmetric
@@
$\mathrm{Singl}\mbox{{AC}}+\mathrm{BLP}$ $\mathrm{CSingl}\mbox{{AC}}+\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette function whose first $n$ blocks are totally symmetric and last block is symmetric.
@@
$\mathrm{Singl}\mbox{{AC}}+\mathrm{AIP}$ $\mathrm{CSingl}\mbox{{AC}}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette function whose first $n$ blocks are totally symmetric and last block is alternating.
@@
Thus, up to domain size 7 (even in the multi-sorted case), all tractable CSPs can be solved by this very simple algorithm. However, this is no longer true for domain size 8: as observed by Kompatscher, the structure on 8 elements arising from the dihedral group $\mathbf{D}_{4}$ does not admit palette symmetric polymorphisms, implying that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ fails to solve it.
@@
As a byproduct, we also show that the two versions of the singleton algorithms — fixing a tuple of a constraint ($\mathrm{CSingl}$) or a value of a variable ($\mathrm{Singl}$) — have the same power for most of the algorithms. Surprisingly, in Section 6 we discovered two structures $\mathbb{A}$ and $\mathbb{B}$, where $\mathbb{A}$ is finite and $\mathbb{B}$ is infinite, such that $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ can be solved by $\mathrm{CSingl}\mathrm{AC}$ but not by $\mathrm{Singl}\mathrm{AC}$. This emphasizes that the equivalence of $\mathrm{Singl}$ and $\mathrm{CSingl}$ for finite structures is nontrivial and the use of the Hales-Jewett theorem is essential.
@@
In Section 4 we define several CSP templates to illustrate the limitations of linear programming algorithms such as BLP, AIP, and $\mathrm{BLP}+\mathrm{AIP}$. By providing concrete unsatisfiable instances on which $\mathrm{BLP}+\mathrm{AIP}$ incorrectly returns “yes”, we demonstrate that the failure of the algorithm arises because the corresponding template is, in a certain sense, disconnected. We then show that the singleton version overcomes this difficulty, as fixing an element allows it to solve each connected part separately. Finally, we show that the template arising from the dihedral group $\mathbf{D}_{4}$ cannot be solved even by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. Thus, we identify another fundamental reason why linear programming algorithms fail. Notice that in this section we not only present instances that fool the algorithms but also demonstrate the characterization of the singleton versions of the algorithms by providing concrete palette symmetric polymorphisms.
In Section 6, we apply our ideas to temporal CSPs, where a relational structure is called temporal if its domain is $\mathbb{Q}$ and its relations are definable by Boolean combinations of atomic formulas of the form $x<y$. The classification of the complexity of temporal CSPs has been known for a long time [9], but only recently Mottet proposed a new uniform algorithm based on universal algorithms [36]. Notice that most of the universal algorithms cannot be applied directly to infinite domain CSPs. For instance, in linear relaxations one introduces a variable for every element of the domain, which would result in a system with infinitely many variables.
One of the ways to overcome this difficulty is “sampling”: to solve an instance with $N$ variables one first restricts the domain to any $N$-element subset of $\mathbb{Q}$, then runs a universal algorithm on the restricted instance, and returns its answer. Mottet proved in [36] that every tractable temporal CSP can be solved this way using local consistency and $\mathrm{AIP}$ algorithms. The soundness of the algorithm follows from the fact that $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ is solvable by the corresponding universal algorithms, where $\mathbb{A}|_{[N]}$ is the induced substructure obtained by restricting the domain to $\{1,2,\dots,N\}$.
In the paper we present an alternative proof of this result by providing concrete palette symmetric functions from $\operatorname{Pol}(\mathbb{A}|_{[N]},\mathbb{A})$, which can be used to obtain a (true) solution from the outputs of the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{CSingl}\mathrm{AIP}$. Thus, we prove the following slight strengthening of Mottet’s result, where local consistency is replaced by $\mathrm{CSingl}\mathrm{AC}$:
Theorem 1.3.
Let $\mathbb{A}$ be a temporal structure, $N\in\mathbb{N}$. Then one of the following holds:
  • $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ is solvable by $\mathrm{CSingl}\mathrm{AC}\wedge\mathrm{Singl}\mathrm{AIP}$;
  • $\operatorname{CSP}(\mathbb{A})$ is NP-hard.